OpenStudy (anonymous):

If x≠-1,x≠1 or x≠3, then what is (x^2-x)/(2x-6)*(x^2-2x-3)/(x^2-1) ?

hartnn (hartnn):

can u factor x^2-2x-3 ?

OpenStudy (anonymous):

yeah (x+1)(x-3)

hartnn (hartnn):

so, x^2-x=x(x-1) x^2-1=(x+1)(x-1) now just substitute these factors and see what cancels.